OpenStudy (anonymous):

The diameter of a circle is 15cm. What is the radius of the circle?

OpenStudy (mertsj):



OpenStudy (mertsj):

That is the diameter

OpenStudy (mertsj):



OpenStudy (mertsj):

That is the radius

OpenStudy (anonymous):

The radius is known to be 1/2 the diameter: \[D=2r\]\[\frac{D}{2}=r\]\[15=2r\]\[r=7.5cm\]

OpenStudy (mertsj):

As you can see, you have to divide the diameter in half to get the radius.

OpenStudy (anonymous):

if 15 is the diameter.... 15 / 2 = 7.5 (radius)

OpenStudy (anonymous):

Right, thank you all.